破壊 · 構造力学特論 2 単位 (選択)

Advanced Fracture and Structural Mechanics  成行義文 · 教授/ 知的力学システム工学専攻建設創造システム工学コース建設構造工学講座

【授業目的】平面骨組構造物のマトリックス変位法による解析方法を理解させ，

例題の解説と演習問題によりその応用力を養う．【授業概要】平面骨組構造物をコンピュータで解析するのに適したマトリックス

変位法について解説し，演習問題を解かせて応用力を養成する．なお，教科書には，英文テキストを使用する．また，輪講形式の授業である．

【授業形式】講義【キーワード】骨組構造，マトリックス変位法，平面構造，弾性解析，英文【先行科目】『構造力学 1』 (1.0)，『構造力学 3』 (1.0)，『応用構造力学』 (0.8)，『構

造解析学及び演習』 (0.8) 【関連科目】『建築構造特論』 (0.5)，『建設創造システム工学論文輪講』 (0.5) 【履修要件】構造力学の知識を有すること．【履修上の注意】授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時

間の復習をしたうえで授業を受けることが，授業の理解と単位修得のために必要である．またレポート提出を忘れないこと．

【到達目標】平面骨組構造物のマトリックス変位法による解析方法を理解する．【授業計画】

1. ガイダンス/構造解析法の変遷 (pp.1-10)

2. 構造解析に関する基礎事項 1(pp.11-16)

3. 構造解析に関する基礎事項 2/静定構造物と不静定構造物 (pp.16-21)

4. 構造解析法 (pp.21-26)

5. 変位法/軸方向力を受ける棒要素の剛性行列 (pp.26-33)

6. 棒構造物の剛性行列 (pp.33-39)

7. 剛性行列の特性/ねじりを受ける棒要素の剛性行列 (pp.39-44)

8. はり要素の剛性行列 (pp.44-47)

9. 直接剛性法による剛性行列の組立て 1(pp.47-54)

10. 直接剛性法による剛性行列の組立て 1(pp.54-59)

11. 対称幾何学 (pp.59-64)

12. 所定の変位に関する 2 ・ 3 の考察 (pp.64-70)

13. 演習問題 (pp.71-72)

14. 演習問題 (pp.72-73)

15. 期末試験

16. 答案返却及び解説【成績評価基準】期末試験の評点 (40 点満点) と和訳レポートの評点 (60 点満点) の

合計が 60 点以上の場合を合格とする．【教科書】 Matrix and finite element displacement analysis of structures, D.J.DAWE,

Clarendon press, Oxford, 1984 【参考書】授業の際に適宜紹介する. 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216812 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 成行 (A510, 088-656-7326, nariyuki@ce.tokushima-u.ac.jp) ���� (オフィスアワー: 年度ごとに学科の掲示を参照すること)

Advanced Fracture and Structural Mechanics 2 units (selection)

Yoshifumi Nariyuki · Professor / Structural Engineering, Civil and Environmental Engineering, Intelligent Structures and Mechanics Systems Engineering

Target⟩ To understand method of matrix displacement analysis of plane framed

structures which is widely used as a suitable method for computer programming. Outline⟩ First, method of matrix displacement analysis of plane framed structurs

is explained. Next, some problems are given for exercises in analysis of them.

Finally, term examination is set. This class is a seminar type using textbook written in English.

Style⟩ Lecture �������⟩ framed structure, matrix displacement method, plane frame, elastic

analysis, English textbook Fundamental Lecture⟩ “Structural Mechanics 1”(1.0), “Structural Mechanics 3”

(1.0), “Applied Structural Mechanics”(0.8), “Structural Analysis with Exercise”

(0.8)

Relational Lecture⟩ “Advanced building construction”(0.5), “Advanced Civil and

Environmental Engineering Seminar”(0.5) Requirement⟩ Students are required to have a good understanding of undergraduate-

level structure mechanics. Notice⟩ Students are requested to do two hours preparation and two hours review

for each lesson. Goal⟩ To understand method of matrix displacement analysis of plane frames

Schedule⟩

1. Guidance/A few historical remarks(pp.1-10)

2. Basic considerations of structural analysis 1(pp.11-16)

3. Basic considerations of structural analysis 2/Determinate and indeterminate structures(pp.16-21)

4. Methods of analysis (pp.21-26)

5. Displacement method/Stiffness matrix of a bar element subjected to axial force(pp.26-33)

6. Bar structure stiffness matrix(pp.33-39)

7. Some properties of stiffness matrices/Stiffness matrix of a bar element subjected to torsion(pp.39-44)

8. Stiffness matrix of a beam element (pp.44-47)

9. Assembly of the structure stiffness matrix by the direct stiffness method 1(pp.47-54)

10. Assembly of the structure stiffness matrix by the direct stiffness method 2(pp.54-59)

11. Symmetrical geometry(pp.59-64)

12. Further remarks on prescribed displacements(pp.64-70)

13. Problem exercises(pp.71-72)

14. Problem exercises(pp.72-73)

15. Term examination

16. Restoration of answer papers and comments

Evaluation Criteria⟩ Term examination and report are marked out of 60 and 40

respectively and those marks are summed up. The passing mark is 60. Textbook⟩ Matrix and finite element displacement analysis of structures, D.J.

DAWE, Clarendon press, Oxford, 1984 Reference⟩ To be introduced in the class

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216812

Student⟩ Able to be taken by only specified class(es)

Contact⟩

⇒ Nariyuki (A510, +81-88-656-7326, nariyuki@ce.tokushima-u.ac.jp) ���� (Office Hour: 年度ごとに学科の掲示を参照すること)